OpenStudy (anonymous):

what is a factored form of 16x^2 -1

OpenStudy (auctoratrox):

(4x+1)(4x-1)

Parth (parthkohli):

\(1 = 1^2, right?\) So, \[\Large \color{purple}{\rightarrow 16x^2 - 1^2 }\] \[\Large \color{purple}{\rightarrow (4x)^2 - 1^2 }\] \[\Large \color{purple}{\rightarrow (4x + 1)(4x - 1) }\]

Parth (parthkohli):

Using the law \(a^2 - b^2 = (a + b)(a - b)\)

OpenStudy (anonymous):

The anwsers are more complex then that idk how to do their version of it...

OpenStudy (auctoratrox):

whats the whole question?

OpenStudy (anonymous):

What is the factored form of 8x^3 + 1 sorry i typed in the wrong question......my bad.

OpenStudy (auctoratrox):

that's a difference of cubes, so we get a^3 + b^3 = (a + b)(a^2 – ab + b^2)

OpenStudy (anonymous):

(2x+1)(4x^2-2x+1)

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow 8x^3 + 1^3 }\] \[\Large \color{purple}{\rightarrow (2x)^3 + 1^3 }\] \[\Large \color{purple}{\rightarrow (2x+ 1)(2x - 2x + 1) }\]

OpenStudy (auctoratrox):

(2x+1)(4x^2 -2x+1)

Parth (parthkohli):

Oh yes, my bad.

OpenStudy (anonymous):

Thanks you guys it was alot of help!